Simplify the following:
(-u2+4u-8) + (-u – 7) - (-8u2 + 9u + 1)
Please help me find the answer
NOTE: all of the 2's are the powers of the number in front of them

Found 2 solutions by checkley71, AnlytcPhil:
Answer by checkley71(8403)

-U~2+4U-8+-U-7+8U~2-9U-1 OR 7U~2-6U-16 OR (7U-2)(U+8)=0 OR 7U-2=0 OR 7U=2 OR U=2/7 & U+8=0 OR U=-8

Answer by AnlytcPhil(1806)

 (-u² + 4u - 8) + (-u – 7) - (-8u² + 9u + 1)

To remove the parentheses put 1's just before them all

1(-u² + 4u - 8) + 1(-u – 7) - 1(-8u² + 9u + 1)

Then use the distributive law remembering to take the sign
+ or - along with 1 when multiplying through:

       -u² + 4u - 8 - u - 7 + 8u² - 9u - 1

     Combine the -u² with the +8u² and get 7u²
     Combine the +4u, -u, and -9u as -6u
     Combine the -8, the -7, and the -1 as -16

                   7u² - 6u - 16      

 If you like, you can factor that as

                  (7u + 8)(u - 2)

Edwin
AnlytcPhil@aol.com